Question 11 The standard cost of product B manufactured by Bhaskara Company includes 3 units of direct materials at $6.55 perDropdown options for all are favorable, unfavorable or neither
Answer #1

(a).

Total Material Variance

= Standard Cost of Production - Actual Cost of Production

= 9500*3*6*55 - 29000*6.16

= 186675 - 178640

= $8035 Favorable

Material Price Variance

= Standard price of Actual quantity used - Actual price of Actual quantity used

= 6.55*29000 - 6.16*29000

= 189950 - 178640

= $ 11310 Favorable

Material Quantity Variance

= Standard quantity of Material for Actual Production at standard price - Actual quantiy of Material for Actual at standard price Production

= = 9500*3*6*55 - 29000*6.55

= 186675- 189950

=$3275 Unfavorable

(b).

Total Material Variance

= Standard Cost of Production - Actual Cost of Production

= 9500*3*6*55 - 28000*6.75

=186675 - 189,000

=$2325 unFavorable

Material Price Variance

= Standard price of Actual quantity used - Actual price of Actual quantity used

=28000*6.55 - 28000*6.75

=$5600 Unfavorable

Material Quantity Variance

= Standard quantity of Material for Actual Production at standard price - Actual quantity of Material for Actual at standard price Production

=9500*3*6*55 – 28000*6.55

=186675- 183400

=$3275 Favorable